# The MUSE eXtremely Deep Field: Detections of circumgalactic Siāii*Ā emission at ā ā thanks: Based on observations made with ESO telescope at the La Silla Paranal Observatory under the large program 1101.A-0127

### v1

@@ Acknowledgements.
- Skidmore etĀ al. (2015) Skidmore, W., DellāAntonio, I., Fukugawa, M., etĀ al. 2015, Thirty Meter Telescope Detailed Science Case: 2015
- Steidel etĀ al. (2011) Steidel, C.Ā C., BogosavljeviÄ, M., Shapley, A.Ā E., etĀ al. 2011, The Astrophysical Journal, 736, 160
- Steidel etĀ al. (2018) Steidel, C.Ā C., BogosavljeviÄ, M., Shapley, A.Ā E., etĀ al. 2018, The Astrophysical Journal, 869, 123
- Steidel etĀ al. (2010) Steidel, C.Ā C., Erb, D.Ā K., Shapley, A.Ā E., etĀ al. 2010, The Astrophysical Journal, 717, 289
- Steidel etĀ al. (2016) Steidel, C.Ā C., Strom, A.Ā L., Pettini, M., etĀ al. 2016, The Astrophysical Journal, 826, 159
- Sugahara etĀ al. (2017) Sugahara, Y., Ouchi, M., Lin, L., etĀ al. 2017, The Astrophysical Journal, 850, 51
- The Astropy Collaboration etĀ al. (2018) The Astropy Collaboration, Price-Whelan, A.Ā M., SipÅcz, B.Ā M., etĀ al. 2018, The Astronomical Journal, 156, 123
- The Astropy Collaboration etĀ al. (2013) The Astropy Collaboration, Robitaille, T.Ā P., Tollerud, E.Ā J., etĀ al. 2013, Astronomy & Astrophysics, 558, A33
- Tumlinson etĀ al. (2017) Tumlinson, J., Peeples, M.Ā S., & Werk, J.Ā K. 2017, Annual Review of Astronomy and Astrophysics, 55, 389, arXiv:1709.09180 [astro-ph]
- Turner etĀ al. (2014) Turner, M.Ā L., Schaye, J., Steidel, C.Ā C., Rudie, G.Ā C., & Strom, A.Ā L. 2014, Monthly Notices of the Royal Astronomical Society, 445, 794
- Tytler etĀ al. (2009) Tytler, D., Gleed, M., Melis, C., etĀ al. 2009
- UrbanoĀ Stawinski etĀ al. (2023) UrbanoĀ Stawinski, S.Ā M., Rubin, K. H.Ā R., Prochaska, J.Ā X., etĀ al. 2023, The Astrophysical Journal, 951, 135
- Urrutia etĀ al. (2019) Urrutia, T., Wisotzki, L., Kerutt, J., etĀ al. 2019, Astronomy & Astrophysics, 624, A141
- Vasan G.Ā C. etĀ al. (2024) Vasan G.Ā C., K., Jones, T., Shajib, A.Ā J., etĀ al. 2024, Spatially Resolved Galactic Winds at Cosmic Noon: Outflow Kinematics and Mass Loading in a Lensed Star-Forming Galaxy at $z=1.87$, arXiv:2402.00942 [astro-ph]
- Verhamme etĀ al. (2018) Verhamme, A., Garel, T., Ventou, E., etĀ al. 2018, Monthly Notices of the Royal Astronomical Society: Letters, 478, L60
- Virtanen etĀ al. (2020) Virtanen, P., Gommers, R., Oliphant, T.Ā E., etĀ al. 2020, Nature Methods, 17, 261
- Wang etĀ al. (2020) Wang, B., Heckman, T.Ā M., Zhu, G., & Norman, C.Ā A. 2020, The Astrophysical Journal, 894, 149
- Werk etĀ al. (2016) Werk, J.Ā K., Prochaska, J.Ā X., Cantalupo, S., etĀ al. 2016, The Astrophysical Journal, 833, 54
- Wisotzki etĀ al. (2016) Wisotzki, L., Bacon, R., Blaizot, J., etĀ al. 2016, Astronomy & Astrophysics, 587, A98
- Wisotzki etĀ al. (2018) Wisotzki, L., Bacon, R., Brinchmann, J., etĀ al. 2018, Nature, 562, 229
- Wolfe etĀ al. (1986) Wolfe, A.Ā M., Turnshek, D.Ā A., Smith, H.Ā E., & Cohen, R.Ā D. 1986, The Astrophysical Journal Supplement Series, 61, 249
- Xu etĀ al. (2023) Xu, Y., Ouchi, M., Nakajima, K., etĀ al. 2023, Stellar and AGN Feedback Probed with Outflows in JWST Galaxies at z=3-9: Implications of Frequent Nearly-Spherical Galactic Fountains, publication Title: arXiv e-prints ADS Bibcode: 2023arXiv231006614X
- Yuma etĀ al. (2017) Yuma, S., Ouchi, M., Drake, A.Ā B., etĀ al. 2017, The Astrophysical Journal, 841, 93
- Yuma etĀ al. (2013) Yuma, S., Ouchi, M., Drake, A.Ā B., etĀ al. 2013, The Astrophysical Journal, 779, 53
- Zabl et al. (2021) Zabl, J., Bouché, N. F., Wisotzki, L., et al. 2021, Monthly Notices of the Royal Astronomical Society, 507, 4294, arXiv:2105.14090 [astro-ph]

### v2

@@ Acknowledgements.
- Skidmore et al. (2015) Skidmore, W., Dell’Antonio, I., Fukugawa, M., et al. 2015, Thirty Meter Telescope Detailed Science Case: 2015
- Steidel et al. (2011) Steidel, C. C., Bogosavljević, M., Shapley, A. E., et al. 2011, ApJ, 736, 160
- Steidel et al. (2018) Steidel, C. C., Bogosavljević, M., Shapley, A. E., et al. 2018, ApJ, 869, 123
- Steidel et al. (2010) Steidel, C. C., Erb, D. K., Shapley, A. E., et al. 2010, ApJ, 717, 289
- Steidel et al. (2016) Steidel, C. C., Strom, A. L., Pettini, M., et al. 2016, ApJ, 826, 159
- Sugahara et al. (2017) Sugahara, Y., Ouchi, M., Lin, L., et al. 2017,ApJ, 850, 51
- The Astropy Collaboration et al. (2018) The Astropy Collaboration, Price-Whelan, A. M., Sipőcz, B. M., et al. 2018, ApJ, 156, 123
- The Astropy Collaboration et al. (2013) The Astropy Collaboration, Robitaille, T. P., Tollerud, E. J., et al. 2013, A&A, 558, A33
- Tumlinson et al. (2017) Tumlinson, J., Peeples, M. S., & Werk, J. K. 2017, ARA&A, 55, 389
- Turner et al. (2014) Turner, M. L., Schaye, J., Steidel, C. C., Rudie, G. C., & Strom, A. L. 2014, MNRAS, 445, 794
- Tytler et al. (2009) Tytler, D., Gleed, M., Melis, C., et al. 2009
- Urbano Stawinski et al. (2023) Urbano Stawinski, S. M., Rubin, K. H. R., Prochaska, J. X., et al. 2023, ApJ, 951, 135
- Urrutia et al. (2019) Urrutia, T., Wisotzki, L., Kerutt, J., et al. 2019, A&A, 624, A141
- Verhamme et al. (2018) Verhamme, A., Garel, T., Ventou, E., et al. 2018, MNRAS: Letters, 478, L60
- Virtanen et al. (2020) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020, Nature Methods, 17, 261
- Wang et al. (2020) Wang, B., Heckman, T. M., Zhu, G., & Norman, C. A. 2020, ApJ, 894, 149
- Werk et al. (2016) Werk, J. K., Prochaska, J. X., Cantalupo, S., et al. 2016, ApJ, 833, 54
- Werk et al. (2014) Werk, J. K., Prochaska, J. X., Tumlinson, J., et al. 2014, ApJ, 792, 8
- Wisotzki et al. (2016) Wisotzki, L., Bacon, R., Blaizot, J., et al. 2016, A&A, 587, A98
- Wisotzki et al. (2018) Wisotzki, L., Bacon, R., Brinchmann, J., et al. 2018, Nature, 562, 229
- Wolfe et al. (1986) Wolfe, A. M., Turnshek, D. A., Smith, H. E., & Cohen, R. D. 1986, The ApJS, 61, 249
- Xu et al. (2023) Xu, Y., Ouchi, M., Nakajima, K., et al. 2023, arXiv eprints [arXiv:2310.06614]
- Yuma et al. (2017) Yuma, S., Ouchi, M., Drake, A. B., et al. 2017,ApJ, 841, 93
- Yuma et al. (2013) Yuma, S., Ouchi, M., Drake, A. B., et al. 2013,ApJ, 779, 53
- Zabl et al. (2021) Zabl, J., Bouché, N. F., Wisotzki, L., et al. 2021, MNRAS, 507, 4294, arXiv:2105.14090 [astro-ph]
- Zhu & Ménard (2013) Zhu, G. & Ménard, B. 2013, ApJ, 770, 130
Appendix A Overview of the sample
The basic information about our sample is summarized in 1. Among the 39 sources, only three sources are selected from udf-10 with an integration time of 31 hours (ID, , and ), and the remaining 36 are from mxdf with integration times ranging from 32 hours to 141 hours. The small number of udf-10 sources is due to the overlaps between mxdf field with udf-10 (see figure 2 in Bacon et al. 2023). The mxdf data are assigned to most of the sources observed both in mxdf and udf-10 because of higher qualities (except for ID in our sample).
Figure 9 shows the distribution of the SFRs and the of our sample, which are derived from spectral energy distribution (SED) fitting in the DR2 catalog (using Prospector, see section 6.4 more details Bacon et al. 2023). Although the 1 uncertainties in SFR and are large, the 39 sources (black circles) lie along the relation for galaxies at – in the same catalog (gray dots, star-formation main sequence). This suggests that our sample is not biased toward starburst galaxies. The 14 sources used in the stacking analysis (black circles without gray squares with error bars) tend to have higher SFR and than the other sources (black circles without enclosed by gray squares), which are expected from the selection criteria with UV magnitudes for the stacked sample.
Next, we examine potential differences between the two fields. Sources in mxdf are widely distributed and include low-mass sources with faint magnitudes (see also Fig. 2). Among the three sources in udf-10 (yellow circles), two are located at the massive end of the sequence in Fig. 9, while the remaining one has a low stellar mass. As expected from the depth of the MUSE data, the fractions of low-mass sources with low SFRs and faint magnitudes in mxdf are higher than those in udf-10, regardless of thresholds. However, because of the small sample size in udf-10 (), our main results are unlikely to be significantly affected by the depth of udf-10. Studying the fractions and extents of halos is beyond the scope of this paper, which requires a larger sample and careful assessment of the inhomogeneous data set (see Appendix B.1 for the examination of the SB limits for our sources). In Appendix B.2, we discuss potential differences in SFR and for individually-detected halos and non-halos.
ID | RA | DEC | [mag] | Integration [h] | Stacked | Individual Halo | |
24+ | 53.160088 | -27.776356 | 2.543301 | 24.0 | 47 | Yes | Yes |
35+,∗ | 53.160587 | -27.776120 | 2.542948 | 24.8 | 31 | Yes | Yes |
50 | 53.162849 | -27.771626 | 3.323586 | 25.4 | 31 | Yes | Yes |
51 | 53.165183 | -27.781614 | 2.228735 | 25.4 | 140 | Yes | Yes |
76 | 53.157832 | -27.777562 | 2.775538 | 26.0 | 48 | Yes | No |
99 | 53.163077 | -27.785605 | 2.543808 | 26.5 | 140 | No | No |
106 | 53.163726 | -27.779076 | 3.276452 | 26.3 | 140 | No | No |
118 | 53.157088 | -27.780269 | 3.017429 | 26.4 | 110 | No | No |
149 | 53.167866 | -27.778617 | 3.718381 | 26.9 | 140 | No | No |
231 | 53.162101 | -27.772563 | 2.446336 | 27.4 | 31 | No | No |
263 | 53.162688 | -27.776638 | 3.1873 | 27.9 | 96 | No | No |
1141 | 53.165074 | -27.793661 | 2.344985 | 25.3 | 113 | Yes | Yes |
1316 | 53.169306 | -27.793799 | 2.575916 | 25.7 | 57 | Yes | No |
1395 | 53.174057 | -27.790421 | 2.478122 | 25.9 | 42 | Yes | No |
1759 | 53.170789 | -27.793294 | 2.676148 | 26.7 | 51 | No | No |
1817 | 53.156111 | -27.789788 | 3.413466 | 26.8 | 95 | No | No |
6298 | 53.169249 | -27.781255 | 3.133906 | 27.6 | 140 | No | No |
6664 | 53.162345 | -27.784443 | 2.392062 | 25.7 | 140 | Yes | No |
6666 | 53.159576 | -27.776719 | 3.43494 | 26.0 | 53 | Yes | No |
6674 | 53.166556 | -27.775259 | 2.543293 | 26.7 | 34 | No | No |
6700 | 53.168281 | -27.781056 | 2.995106 | 25.4 | 140 | Yes | No |
6999 | 53.159601 | -27.791271 | 3.427612 | 28.0 | 140 | No | No |
7599 | 53.154801 | -27.780390 | 2.586264 | 27.1 | 32 | No | No |
7600 | 53.153765 | -27.784885 | 2.570747 | 25.2 | 55 | Yes | No |
7615 | 53.170752 | -27.785994 | 2.582759 | 26.9 | 140 | No | No |
7627 | 53.170800 | -27.783148 | 2.86605 | 26.2 | 140 | No | No |
7697 | 53.169279 | -27.781394 | 2.777275 | 27.8 | 140 | No | No |
8048 | 53.167939 | -27.792722 | 2.842856 | 25.3 | 133 | Yes | No |
8277 | 53.167383 | -27.790112 | 2.764135 | 26.7 | 140 | No | No |
8314 | 53.174925 | -27.785724 | 2.69058 | 28.0 | 75 | No | No |
8336 | 53.157657 | -27.783316 | 2.642334 | 27.0 | 140 | No | No |
8378 | 53.159007 | -27.781481 | 3.081696 | 28.0 | 140 | No | No |
8414 | 53.171088 | -27.789138 | 2.844166 | 27.6 | 140 | No | No |
8471 | 53.169458 | -27.778191 | 2.806879 | 26.5 | 121 | No | No |
8501 | 53.167806 | -27.791894 | 2.841994 | 26.4 | 140 | No | No |
8502 | 53.174309 | -27.787684 | 2.07006 | 26.2 | 86 | No | No |
8507 | 53.162236 | -27.783618 | 3.760994 | 27.8 | 140 | No | No |
8513 | 53.158786 | -27.784608 | 2.162213 | 25.9 | 140 | Yes | No |
8537 | 53.169650 | -27.788081 | 3.189507 | 26.1 | 140 | No | No |
Appendix B Potential effects of different depths and sample properties
B.1 Potential effects of different depths of the MUSE data for our sample
As mentioned in Section 3.2, we have five individual detections of Si ii* 1533 halos. Three of these sources are located in udf-10 with the integration time of 31 hours, while the remaining two are in mxdf with the integration times ranging from 32 hours to 141 hours (see Fig. 2 and Table 1). The SB limits of the NBs used in this work depend not only on the integration time but also on the observed wavelength (i.e., redshifts of the sample). The effect of cosmic dimming is also significant and cannot be ignored. Therefore, we discuss potential effects of 1 SB limits and cosmic dimmings on the detectability. If the origin of Si ii* emission is continuum pumping as suggested in this paper, bright UV magnitudes are necessary to produce bright Si ii* emission (in terms of observed, apparent magnitudes). We examine distributions of 1 SB limits on a 1-arcsec2 scale for the Si ii* 1533 NBs and apparent UV magnitudes for our sources in Fig. 10 (with integration times indicated by the color bar). We estimate the 1 SB limits for each object from a sensitivity table for mxdf and udf-10 in Bacon et al. (2023), considering the narrow-band widths and integration times (see the left panel). The right panel shows the 1 SB limits multiplied by (1+)4/(1+)4 to account for the effect of cosmic dimming in a relative manner, where is the systemic redshift of our best-case object, ID (). For instance, the actual depth of the Si ii* NB of ID is erg s-1 cm-2 arcsec-2 (1, left panel). Additionally, because of the cosmic dimming effect, the SB level of ID at becomes times fainter compared to the case where it was located at . In relative terms, it corresponds to the 1 SB limit of erg s-1 cm-2 arcsec-2 (right panel). Considering these two effects, the corresponding noise level of ID is seven times higher than that of ID, which is presented relatively as a correction in the right panel. The noise levels shown in the right panel are less dependent on the integration time (different fields) than those in the left panel. This exercise indicates that the depth of data indeed depends on the redshifts as well as integration times.
As shown in Fig. 10, we cannot conclude any dependence of Si ii* halo detections (red crosses) and nondetections on the SB limits as well as integration times. Even when focusing on bright objects with , which are brighter than the faintest individually-detected halo (ID), and whose IDs are shown in the left panel, no clear trends are observed. The SB limits of two non-halos (ID= at and ID= at in mxdf) are deeper than that of ID= at in udf-10. We note that the Si ii* 1533 of one of the UV brightest objects, ID=, drops in the AO gap in the MXDF and is not included in this plot. A larger sample is needed to rigorously test the detectability depending on the depth of the dataset. In the next section, we discuss differences in sample properties.
B.2 Potential differences in properties between Si ii*-halo sources and non-halo sources
We check potential differences in SFR and between Si ii*-halo sources (red crosses) and non-halo sources (black circles) shown in Fig. 9, though their uncertainties are large. We find that halos tend to have high SFR and , which is expected from the trend with UV magnitudes (Figs. 2 and 10). However, not all the sources with high SFR, high , or high () show a clear Si ii* halo. Among them, apparent UV magnitudes (i.e., and redshifts) would be most strongly related to the detection of the halos. It is indeed consistent with the continuum pumping as the mechanism of Si ii* emission. We need a larger sample to investigate the dependence of the detectability and existence of halos on data qualities and sample properties.